OpenStudy (anonymous):

.In a study of 225 adults, the mean heart rate was 72 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 10 beats per minute. What is the 90% confidence interval for the mean beats per minute? 70.7 – 73.3 70.9 – 73.1 70.7 – 73.1 70.9 – 73.3

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

I am not good at this.. :(

OpenStudy (kropot72):

The confidence interval for the mean is: \[\large (\bar {x}-z\frac{\sigma}{\sqrt{n}}, \bar{x}+z\frac{\sigma}{\sqrt{n}})\] where z is 1.645 to find a 90% confidence interval. Plugging in the given values we get: \[\large CI=([72-1.645\frac{10}{\sqrt{225}}],\ [72+1.645\frac{10}{\sqrt{225}}])\] which simplifies to: \[\large CI=([72-1.1], [72+1.1])\] You can do the final arithmetic.